One end of a uniform 4.30-m-long rod of weight Fg is supported by a cable at an angle of ? = 37
Answer #1

L = length of the rod = 4.3 m
Fg = weight of the rod = weight of the ball
? = angle of the cable = 37
Answer #2

along horzantal

N - T*cos? = 0

N = T*cos?


frictional force f = u*N = u*T*cos?

along vertical

Fnety = 0


f + T*sin? = Fg + Fg = 2Fg


u*T*cos? + T*sin? = 2Fg


u*T*cos? + T*sin? = 2Fg


T = Fg / (u*cos? + sin?)

net torque = 0

Fg*(L/2) + w*x = T*sin? *L


(L/2) + x = sin?*L / (u*cos? + sin?)

x = sin?*L / (u*cos? + sin?) - L/2


x = Fg*sin37*4.3 / (0.54*cos37 + sin37) - 2.15


2.15 + x = 2.505


x = 0.355m = 35.5 cm
Answer #3

along horzantal

N - T*cos? = 0

N = T*cos?


frictional force f = u*N = u*T*cos?

along vertical

Fnety = 0


f + T*sin? = Fg + Fg = 2Fg


u*T*cos? + T*sin? = 2Fg


u*T*cos? + T*sin? = 2Fg


T = Fg / (u*cos? + sin?)

net torque = 0

Fg*(L/2) + Fg*x = T*sin? *L = Fg*sin?*L / (u*cos? + sin?)


(L/2) + x = sin?*L / (u*cos? + sin?)

x = sin?*L / (u*cos? + sin?) - L/2


x = sin37*4.3 / (0.545*cos37 + sin37) - 2.15


2.15 + x = 2.505


x = 0.345m = 34.5 cm